One charged plastic ball is vertically above anoth
Answer #1
2d
Answer #2
2d
Answer #3
mg = kq/d*d

mg = k 4q*q/x*x

equating them

1/d*d = 4/x*x

x = 2d answer
Answer #4

the charge down is basically balancing the gravitational force of the upper ball to keep it in equlibrium ...

if the charges are doubled then kq1*q2 becomes 4 times the initial value to compensate this if the distance between the charges has to be also doubled ...(as weight of the ball remains constant during injection of charge )

therefore the distance between them should also be doubled

2d
Answer #5
8d